# Mixture of Directed Graphical Models for Discrete Spatial Random Fields

### v3

@@ 3.2 MRF Prior
As stated in the introduction, MRFs are popular because of the intuitive representation of the probability distribution through the NUG. In particular, undirected graphs satisfy the local Markov property which states that a vertex, $i$, is independent of all other vertices given its neighbors (See Appendix C for a more thorough discussion). The Markov blanket of a vertex $i$ is the set of neighbor vertices and as such the full conditionals $p(z_{i}|\mathbf{z}_{-i},\boldsymbol{\xi})$ from an MRF specification reduce to $p(z_{i}|\mathbf{z}_{\partial(i)},\boldsymbol{\xi})$. While an MRF is defined as a probability distribution whose conditional distributions define a NUG (Cressie, 1993, pg. 415), we can, in actuality, start with a collection of full conditional distributions consistent with a NUG, $p(z_{i}|\mathbf{z}_{\partial(i)},\boldsymbol{\xi})$, and obtain the corresponding MRF, which we are guaranteed is a valid joint probability distribution by the Hammersley-Clifford theorem (an extended proof given by Besag (1974))
The result depends on the notion of a clique. Any single node or set of vertices which are all mutually neighbors are defined to be a clique. Let $C=\{c_{1},\dots,c_{k}\}$ be the set of all cliques in the graph, $\mathbf{N}$. From the Hammersley-Clifford Theorem, if the probability distribution of a MRF, $p(\mathbf{z}|\boldsymbol{\xi})$ for $\mathbf{z}\in\mathcal{Z}$, with respect to graph $\mathbf{N}$ and cliques $C$, satisfies $p(\mathbf{z}|\boldsymbol{\xi})>0$ for all $\mathbf{z}\in\mathcal{Z}$, then we can write
@@
Here $\psi_{c}(\mathbf{z}_{c}|\boldsymbol{\xi}_{c})$ are arbitrarily chosen, nonnegative and finite functions parameterized by $\boldsymbol{\xi}=\{\boldsymbol{\xi}_{c}:c\in C\}$ and the partition function,
@@
$p(\mathbf{z}|\beta)\propto\exp\left(\beta\sum_{i\sim j}I(z_{i}=z_{j})\right),$
@@
a version of the Ising model from statistical mechanics as it is commonly parameterized in the spatial statistics setting: $\beta$ governs the number of matches of the values of the vertices with respect to the edges of the graph. As $\beta$ increases, the black and white areal units will appear to be more clustered.
A MRF prior specification for $p(\mathbf{z}|\boldsymbol{\theta})=p(\mathbf{z}|\boldsymbol{\xi})$ also leads to a valid posterior distribution. The difficulty of sampling from the posterior in an MCMC scheme comes from the partition function $R(\boldsymbol{\xi})$, which involves the unknown parameters $\boldsymbol{\xi}$. Methods exist to obtain valid posterior samples by avoiding evaluation of the partition function (Møller et al., 2006; Murray et al., 2006), but rely on the computationally burdensome coupling from the past algorithm of Propp and Wilson (1996).
@@
We now introduce the popular aMRF model where the analytical form of the PL is used in place of an MRF prior to facilitate Bayesian inference via MCMC. In the aMRF we replace the MRF distribution, $p(\mathbf{z}|\boldsymbol{\xi})$, with the tractable approximation
@@
the product of the full conditionals. For our model comparisons, we use the full conditionals of Equation 6 for the aMRF prior. Besag (1975) named $g(\mathbf{z}|\boldsymbol{\xi})$ the psuedo-likelihood and used a factorization technique to show that the PL is a consistent estimator of $\boldsymbol{\xi}$, but notably did not suggest it replace the analytical form of the MRF in a Bayesian setting. Regardless, we often find that inference is then carried out using the approximation
@@
We provide a proof of Proposition 3.1 in Appendix B.1 using the same factorization technique as in Besag (1975). Now, as $g(\mathbf{z}|\boldsymbol{\xi})$ is not a valid probability distribution, there is no guarantee (that we are aware of) that sampling from the right hand side of Equation 7 (as described in Section 4) will yield draws from a valid posterior distribution, let alone the posterior in Equation 2.
@@ 4 Model Fitting
We carry out posterior inference through Markov chain Monte Carlo (MCMC). We present a general MCMC algorithm that can accommodate all three of the spatial priors described above with a generic class of compatible DAGs for the MDGM. We detail modifications for updating the graph in the mixture model for individual classes presented previously. Model fitting is facilitated by treating the DGM as a hidden state which can be updated for each $b=0,\dots,B-1$ iterations of the MCMC.
@@
For our MDGM models, the MCMC algorithm begins by updating the the current DAG. (For an MCMC algorithm for fitting a model with an MRF prior on the latent variable, the NUG is fixed, so this step is skipped.) We initialize the algorithm with starting values for $\boldsymbol{\eta}^{0},\beta^{0},\mathbf{z}^{0}$ and $\mathbf{D}^{0}$. For the AO and rooted classes, we update the DAG using an independent Metropolis-Hastings (MH) step. Let $q(\mathbf{D})$ be the proposal distribution for a new DAG. For both the rooted and AO classes, we set the proposal distribution to be the prior. A uniform draw from rooted class is obtained by randomly selecting a root vertex then generating the $i$-rooted DAG from Algorithm 1. In the case of the AO class, we obtain an acyclic orientation through permutation of the vertices to assign directions to the edges in the graph (Section 2.2.1). Given current DAG, $\mathbf{D}^{b}$, and proposal, $\mathbf{D}^{*}$, set $\mathbf{D}^{b+1}=\mathbf{D}^{*}$ with probability
For the ST class, we can sample $\mathbf{D}\in\mathcal{D}^{ST}(\mathbf{N})$ directly from the full conditional posterior distribution $p(\mathbf{D}|\mathbf{z},\beta,\mathbf{y})$. We can write the posterior as
$p(\mathbf{D}|\mathbf{z},\beta,\mathbf{y})$ $\propto p(\mathbf{z}|\mathbf{D},\beta)p(\mathbf{D})$ (9)
and since $|\pi(i)|=1$, we have
$p(\pi(i)=j|\mathbf{z},\beta,\mathbf{y})\propto\frac{\exp\left(\beta I(z_{i}=z_{j})\right)}{\exp\left(\beta I(z_{j}=0)\right)+\exp\left(\beta I(z_{j}=1)\right)},$
for $j\in\partial(i)$. Thus, a draw of a spanning tree from the posterior can be obtained using Wilson’s algorithm as described in Section 2.2.3 with the modification that the random walk occurs with edge weights given by the above. This modification explicitly generates a ST-DAG with root at the vertex first selected to be part of the spanning tree, yet since any rooted ST-DAG with the same tree structure are equivalent, and therefore nonidentifiable, it does not matter which vertex rooted ST-DAG we choose from a given undirected tree. That is to say, selecting the root vertex is not informed by the data and can be chosen uniformly at random without affecting the posterior probability of the undirected tree. We use the ST-DGM formulation of $p(\mathbf{z}|\mathbf{D},\beta)$ to facilitate unified computation between the MDGM classes.
where $z_{i}$ depends on its the Markov blanket defined by the current DAG, $\mathbf{D}^{b+1}$, and the observed outcome, $\mathbf{y}_{i}=[y_{i1},\dots,y_{im_{i}}]$. For the MRF prior (for both exact inference or the aMRF), the full conditionals obtain an update $\mathbf{z}^{b+1}$ are
Next, we update the spatial dependence parameter, $\beta$, through a MH step. Let $\beta^{*}$ be the proposal drawn from $q(\beta^{*}|\beta^{b})$. Then, set $\beta^{b+1}=\beta^{*}$ with probability
where $f(\mathbf{z}^{b+1}|\beta^{*})$ is a place holder for $p(\mathbf{z}|\mathbf{D},\beta)$, the MDGM prior, and $g(\mathbf{z}|\beta)$ when using an aMRF prior. Exact inference can be carried out with the clever MH ratio of Møller et al. (2006) and Murray et al. (2006), which avoids computing the normalizing constant, though relies on generating an exact sample from $p(\mathbf{z}|\beta^{*})$ using the coupling from the past algorithm (Propp and Wilson, 1996). Generate an exact sample $\mathbf{z}^{*}$ from $p(\mathbf{z}|\beta^{*})$ and let
Then, set $\beta^{b+1}=\beta^{*}$ with probability
Lastly, we update the noise parameter $\boldsymbol{\eta}$. When $p(\eta)\equiv\mathrm{Beta}(\eta|a,b)$ then the full conditionals are both truncated beta distributions,
$p(\eta_{1}|\eta_{0},\mathbf{y},\mathbf{z})\propto\\ \mathrm{Beta}\left(\eta_{1}|a_{1}+\sum_{i=1}^{n}\sum_{j=1}^{m_{i}}y_{ij}I(z_{i}=1),\;b_{1}+\sum_{i=1}^{n}\sum_{j=1}^{m_{i}}(1-y_{ij})I(z_{i}=1)\right)I(\eta_{1}>\eta_{0})$
and
$p(\eta_{0}|\eta_{1},\mathbf{y},\mathbf{z})\propto\\ \mathrm{Beta}\left(\eta_{0}|a_{0}+\sum_{i=1}^{n}\sum_{j=1}^{m_{i}}y_{ij}I(z_{i}=0),\;b_{0}+\sum_{i=1}^{n}\sum_{j=1}^{m_{i}}(1-y_{ij})I(z_{i}=0)\right)I(\eta_{1}>\eta_{0}).$
Table 1 gives a summary of the MCMC algorithms for our MDGM classes and for the standard MRF prior, both for exact inference and for the aMRF.
@@ 5 Model Evaluation
As we propose the MDGM as an alternative for modeling spatial dependence, we compare how well each class of the MDGM performs to the classic MRF, as well as to the aMRF, a persistent default substitute for the MRF in Bayesian analyses. For our simulation study we generate data from our model described in Section 3, but using an MRF as the data generating mechanism for the latent variable. For a single iteration, we generate a latent binary field, $\mathbf{z}$, for a specified NUG and spatial dependence parameter $\beta$. Then, given the latent $\mathbf{z}$, we generate a vector of independent binary observations $\mathbf{y}_{i}$ with Bernoulli errors $\eta_{0}$ and $\eta_{1}$ at each areal unit for $i=1,\dots,n$. Using only the observed $\mathbf{y}$ as input, we perform full Bayesian analysis using the MCMC algorithms described in Section 4 for the ST, rooted, and AO classes of the MDGM as well as for the MRF both without any approximations and using an aMRF. The main goal of the simulation study is to evaluate “how well” each model estimates the latent spatial field given the observed binary response.
@@
We evaluate model performance by the posterior mean accuracy of the estimated latent field to the ground truth. That is, we generate the posterior distribution of accuracy by finding the rate of correspondence between each draw of the latent field and the ground truth and average over these individual estimates to obtain the posterior mean accuracy. Additionally, we evaluate model performance by how well the different models allow us to estimate the true strength of spatial dependence assumed in the data generation process. Using the MRF as the data generating mechanism gives us an inherent metric to evaluate spatial dependence, namely its sufficient statistic $T(\mathbf{z})=\sum_{i\sim j}I(z_{i}=z_{j})$, the number of neighboring pairs of areal units with the same value. For a single iteration of the simulation study, we find the posterior root mean square error (RMSE) of the sufficient statistic of the draws from the posterior to the $T(\mathbf{z})$ of the generated data. Note, we are not interested in evaluating how well each model estimates the spatial dependence parameter, $\beta$, as the mathematical interpretation of this parameter is inherently different for each of the classes of the MDGM and for an MRF. For each distinct setting in our simulation study, i.e. different values of $\beta$, $\boldsymbol{\eta}$, and $m_{i}$ (the number of observations at each areal unit), we generate one hundred datasets and summarize model performance by estimating the expected posterior mean accuracy and expected posterior RMSE of $\hat{T}(\mathbf{z})$ by taking the average across the hundred iterations. For the MCMC algorithms, we initialize the parameters at the true values of $\beta$ and $\boldsymbol{\eta}$ and an independently and randomly generated $\mathbf{z}$. We generate two thousand draws from the posterior, discarding the first one thousand as burn in. The simulation study was performed in R (R Core Team, 2024).
@@
We evaluate the MDGM models a missing data setting, with a fixed rate of missingness across the areal units. We also include a complete data setting, that is at least one observed $y_{i}$ for $i=1,\dots,n$, in Appendix D.2. The latent field is generated for a $16\times 16$ regular lattice (256 areal units) with a second-order dependence structure as this structure admits a rooted class of DAGs which are Markov blanket equivalent with the specified NUG. Additionally, the smaller sized lattice allows us to perform the computationally expensive exact inference for the MRF. We use the exact.mrf function from the GiRaF package (Stoehr et al., 2020) in R to obtain a draw from the MRF distribution needed for the Metropolis-Hastings ratio in Equation 11. For all simulations we set $\eta_{0}=\eta$ and $\eta_{1}=1-\eta$. First, we generate $m_{i}\sim\mathrm{pois}(\lambda)$. We set $\eta=0.1$ and perform the simulation across the combinations of $\beta=\{0.1,0.15,0.2,0.25,0.3\}$ and $\lambda=\{1.39,2.3\}$, which corresponds to a missing rate of 25% and 10% respectively. The selection of the values for the spatial dependence parameter, $\beta$, to use in the simulation study were driven by a well studied property of the Ising model in statistical mechanics and statistics, phase transition (see Georgii, 2011), which we explain in Appendix D.1.
@@
Across all iterations of the simulation study, we found that the average elapsed computation time of the aMRF was the fastest, with roughly two and a half minutes to obtain two thousand draws from the posterior. The MDGM models all relatively had similar average elapsed times ranging from eight minutes for the ST class to ten minutes for the rooted class. The exact MRF model was nearly eight times slower than the MDGM models, with an average elapsed time of 63 minutes.
@@
The results of the simulation study, the estimates of the expected statistic with a 90% bootstrap confidence interval of the Monte Carlo error, are shown in Figure 1. The bootstrap confidence interval estimates were obtained by taking one thousand samples, of size one hundred and with replacement, of the expected posterior statistics across the simulation studies, then taking the average of each resample to estimate the sampling distribution. We took the fifth and ninety fifth quantiles of the estimated sampling distribution to be our the lower and upper bounds of the confidence interval.
@@
From the simulation study two overarching trends emerge. First, in the low uncertainty setting ($\lambda=2.3$, 10% missing rate), all of the models seem to perform equally well, both in terms of accuracy and in estimating the spatial dependence of the latent variable. Second, we observe poorer performance of the AO and rooted classes of the MDGM and aMRF in the higher uncertainty setting. In particular, as the true spatial dependence parameter increases (the latent field becomes more clustered), the aMRF appears to be the worst performing model in estimating the spatial dependence of the latent variable. In comparison, the AO and rooted classes have relatively worse performance in estimating spatial dependence when the true spatial dependence parameter is low, but show an increase in relative performance to the aMRF as $\beta$ increases. Throughout the simulation studies, the exact solution performs the best, which is to be expected as the data were generated from an MRF. Notably in the high uncertainty regime, the MDGM-ST is consistently the next best performing model, particularly in estimating the spatial dependence of the latent variable.

### v4

@@ 3.2 MRF Prior
We now present the MRF prior to which we will compare our new MDGM framework. As stated in the introduction, MRFs are popular because of the intuitive representation of the probability distribution through the NUG. In particular, undirected graphs satisfy the local Markov property which states that a vertex, $i$, is independent of all other vertices given its neighbors. (See Appendix G for a more thorough discussion.) From the local Markov property, the full conditionals, $p(z_{i}|\mathbf{z}_{-i},\boldsymbol{\xi})$, from an MRF specification reduce to $p(z_{i}|\mathbf{z}_{\partial(i)},\boldsymbol{\xi})$. While an MRF is defined as a probability distribution whose conditional distributions define a NUG (Cressie, 1993, pg. 415), we can, in actuality, start with a collection of full conditional distributions consistent with a NUG, $p(z_{i}|\mathbf{z}_{\partial(i)},\boldsymbol{\xi})$, and obtain the corresponding MRF, which we are guaranteed is a valid joint probability distribution by the Hammersley-Clifford Theorem, for which an extended proof given by Besag (1974).
The result depends on the notion of a clique. Any single node or set of vertices which are all mutually neighbors are defined to be a clique. Let $C=\{c_{1},\dots,c_{k}\}$ be the set of all cliques in the graph, $\mathbf{N}$. From the Hammersley-Clifford Theorem, if the probability distribution of an MRF, $p(\mathbf{z}|\boldsymbol{\xi})$ for $\mathbf{z}\in\mathcal{Z}$, with respect to graph $\mathbf{N}$ and cliques $C$, satisfies $p(\mathbf{z}|\boldsymbol{\xi})>0$ for all $\mathbf{z}\in\mathcal{Z}$, then we can write
@@
Here $\psi_{c}(\mathbf{z}_{c}|\boldsymbol{\xi}_{c})$ are arbitrarily chosen, nonnegative, and finite functions parameterized by $\boldsymbol{\xi}=\{\boldsymbol{\xi}_{c}:c\in C\}$ and thepartition function,
@@
$p(\mathbf{z}|\beta)\propto\exp\left(\beta\sum_{i\sim j}I(z_{i}=z_{j})\right),$ (7)
@@
a version of the Ising model from statistical mechanics as it is commonly parameterized in the spatial statistics setting: $\beta$ governs the number of matches of the values of the vertices with respect to the edges of the graph. As $\beta$ increases, the areal units will appear to be more clustered.
An MRF prior specification for $p(\mathbf{z}|\boldsymbol{\theta})=p(\mathbf{z}|\boldsymbol{\xi})$ also leads to a valid posterior distribution. The difficulty of sampling from the posterior in an MCMC scheme comes from the partition function $R(\boldsymbol{\xi})$, which involves the unknown parameters $\boldsymbol{\xi}$. Methods exist to obtain valid posterior samples by avoiding evaluation of the partition function (Møller et al., 2006; Murray et al., 2006), but rely on the computationally burdensome coupling from the past algorithm of Propp and Wilson (1996).
@@
As another prior for comparison to the MDGM, we include the popular aMRF model (our name) where the analytical form of the pseudo-likelihood is used in place of an MRF prior to facilitate Bayesian inference via MCMC. In the aMRF prior, we replace the MRF distribution, $p(\mathbf{z}|\boldsymbol{\xi})$, with the tractable approximation
@@
the product of the full conditionals. For our model comparisons, we use the full conditionals of Equation 8 for the aMRF prior. Besag (1975) named $g(\mathbf{z}|\boldsymbol{\xi})$ the psuedo-likelihood and used a factorization technique to show that the pseudo-likelihood is a consistent estimator of $\boldsymbol{\xi}$, but notably did not suggest it replace the analytical form of the MRF in a Bayesian setting. Regardless, we often find that inference is then carried out using the approximation
@@
We provide a proof of Proposition 3.1 in Appendix F.1 using the same factorization technique as in Besag (1975). Now, as $g(\mathbf{z}|\boldsymbol{\xi})$ is not a valid probability distribution, there is no guarantee (that we are aware of) that sampling from the right hand side of Equation 9 (as described in Section 4) will yield draws from a valid posterior distribution, let alone the posterior in Equation 2.
@@ 4 Model Fitting
We carry out posterior inference through Markov chain Monte Carlo (MCMC). The MCMC scheme for each spatial prior (MDGM, MRF, and aMRF) all follow the same sequence of steps. First in the MDGM models, model fitting is facilitated by treating the DAG as a hidden state which can be updated for each $b=0,\dots,B-1$ iterations of the MCMC. Since the NUG is fixed for the MRF and aMRF, the graph update step is skipped for these priors. Next, we update the configuration of the latent variable, $\mathbf{z}$. Third, we update the spatial dependence parameter, $\beta$, with a Metropolis-Hastings step. In the MRF, this MH step requires an exact sample from the MRF distribution to cleverly avoid calculating the normalizing constant (Møller et al., 2006; Murray et al., 2006). Lastly, given the latent configuration $\mathbf{z}$, the update for $\boldsymbol{\eta}$ is the same for all models. Table 1 gives a summary of the MCMC algorithms for our MDGM classes and for the standard MRF prior, both for exact inference and for the aMRF. We refer the reader to Appendix E detailed executions of each step in the presented algorithms and the details for the time complexity. Of note for the spanning tree class, the run time depends on the structure of the NUG, with denser NUGs enabling faster sampling of spanning trees from the posterior. In the standard spatial setting, the time complexity of MCMC with MDGM-ST prior is $O(n\log n)$. For the models with MDGM-R, MDGM-AO and aMRF priors the time complexity of the MCMC is $O(n)$, whereas MCMC for exact inference with an MRF prior has time complexity $O(n^{3})$. See Appendix E.5 for more details.
@@ 5 Model Evaluation
As we propose the MDGM as an alternative for modeling spatial dependence, we now compare how well two classes of the MDGM perform to the classic MRF, as well as to the aMRF, a persistent default substitute for the MRF in Bayesian analyses. For our simulation study we generate data from our model described in Section 3, but using an MRF as the data generating mechanism for the latent variable. For a single simulated dataset, we generate a latent binary field, $\mathbf{z}$, for a specified NUG and spatial dependence parameter $\beta$. Then, given the latent $\mathbf{z}$, we generate a vector of independent binary observations $\mathbf{y}_{i}$ with Bernoulli errors $\eta_{0}$ and $\eta_{1}$ at each areal unit for $i=1,\dots,n$. Using only the observed $\mathbf{y}$ as input, we perform full Bayesian analysis using the MCMC algorithms described in Section 4 and Appendix E for the spanning tree and acyclic orientation classes of the MDGM as well as for the MRF both without any approximations and using an aMRF. The main goal of the simulation study is to evaluate “how well” each model estimates the latent spatial field given the observed binary response.
@@
We evaluate model performance by the posterior mean accuracy of the estimated latent field to the ground truth. That is, we generate the posterior distribution of accuracy by finding the rate of correspondence between each draw of the latent field and the ground truth and average over these individual estimates to obtain the posterior mean accuracy. Additionally, we evaluate model performance by how well the different models allow us to estimate the true strength of spatial dependence assumed in the data generation process. Using the MRF as the data generating mechanism gives us an inherent metric to evaluate spatial dependence, namely its sufficient statistic $T(\mathbf{z})=\sum_{i\sim j}I(z_{i}=z_{j})$, the number of neighboring pairs of areal units with the same value. For a single dataset of the study, we find the posterior root mean square error (RMSE) of the sufficient statistic of the draws from the posterior to the $T(\mathbf{z})$ of the generated data. Note, we are not interested in evaluating how well each model estimates the spatial dependence parameter, $\beta$, as the mathematical interpretation of this parameter is inherently different for each of the classes of the MDGM and for an MRF.
Again, we emphasize that the MDGM is proposed as an alternative to, not approximation of, an MRF; nevertheless, due to the prevalence of the MRF as a model for spatial dependence between areal units, we also include a measure of how well the posterior of each model with an MDGM or aMRF prior approximates the posterior of the model with an MRF prior when the true data generating mechanism is known to follow an MRF. For each simulated dataset we compare the posterior distribution of $T(\mathbf{z})$ for each MDGM class and the aMRF to the posterior distribution of $T(\mathbf{z})$ from an MRF through the total variation distance. The model with an MRF prior is correctly specified and can be viewed as the true posterior for the simulated data. We calculate the empirical total variation distance from $B$ posterior draws as
$\hat{\delta}(p(T|\mathbf{y}),p^{*}(T|\mathbf{y}))=\frac{1}{2B}\sum_{t=0}^{|\bar{E}|}\sum_{b=1}^{B}\left|I(T(\mathbf{z}^{b})=t)-I(T(\mathbf{z}^{*b})=t)\right|,$
where $p(T|\mathbf{y})$ is the posterior of $T(\mathbf{z})$ with an MRF prior and $p^{*}(T|\mathbf{y})$ is the posterior of $T(\mathbf{z})$ with an alternate prior. Correspondingly, $\mathbf{z}^{b}$ is the $b$th draw using the MRF prior and $\mathbf{z}^{*b}$ is the $b$th draw using the alternative prior.
For each distinct setting in our simulation study, i.e. different values of $\beta$, $\boldsymbol{\eta}$, and $m_{i}$ (the number of observations at each areal unit), we generate one hundred datasets and summarize model performance by estimating the expected posterior mean accuracy, expected posterior RMSE of $\hat{T}(\mathbf{z})$ and expected total variation distance by taking the average across the posteriors fit to the hundred simulated datasets. For the MCMC algorithms, we initialize $\boldsymbol{\eta}^{0}$ at the simulation setting of $\boldsymbol{\eta}$, $\beta^{0}$ at the simulation setting for the MRF and aMRF priors and at $2\beta$ for the MDGM priors, and $\mathbf{z}^{0}$ at a independently and randomly generated configuration. We generate five thousand draws from the posterior, discarding the first one thousand as burn-in. We calculate the split potential scale reduction factor (R-hat) for the posterior draws of $\beta,\eta_{0},\eta_{1}$ and $T(\mathbf{z})$ to assess convergence of each posterior (Gelman et al., 2013). Posteriors that did not converge in one of the parameters or $T(\mathbf{z})$ (R-hat greater than 1.1) were not included in the Monte Carlo estimate. Only the MDGM-AO and aMRF indicated a lack of convergence in the high uncertainty regimes (see below) for about 12 posteriors in each simulation setting when $\beta\in\{0.6,0.7,0.8\}$. The data generation was performed in Julia (Bezanson et al., 2017) and the posterior simulation was performed in R (R Core Team, 2024).
@@
We evaluate the MDGM models in the missing data setting, with a fixed rate of missingness across the areal units. We also include a complete data setting, that is, at least one observed $y_{i}$ for $i=1,\dots,n$, in Appendix H.2. We find that the latent spatial field is only partially identifiable (Gustafson, 2015) when $1\geq m_{i}\geq 0$ for $i=1,\dots,n$. Correspondingly, in all settings of the simulation study, we have $m_{i}>1$ for at least one areal unit, a constraint that improves identifiability of the $\boldsymbol{\eta}$ and $\mathbf{z}$. The latent field is generated for a $16\times 16$ regular lattice (256 areal units) with a first-order dependence structure, a common default. The smaller sized lattice allows us to perform the computationally expensive exact inference for the MRF. In Appendix H, we also evaluate the MDGM-ST and aMRF on a $32\times 32$ regular lattice (1024 areal units) in the missing and complete data settings under the high uncertainty regimes. We use the exact.mrf function from the GiRaF package (Stoehr et al., 2020) in R to obtain a draw from the MRF distribution needed for the Metropolis-Hastings ratio in Equation 13. For all simulated datasets we set $\eta_{0}=\eta$ and $\eta_{1}=1-\eta$. First, we generate $m_{i}\sim\mathrm{pois}(\lambda)$. We set $\eta=0.1$ and generate one hundred datasets across each combination of $\beta=\{0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8\}$ and $\lambda=\{1.39,2.3\}$, which corresponds to a missing rate of 25% and 10% respectively. The selection of the values for the spatial dependence parameter, $\beta$, to use in the simulation study were driven by a well studied property of the Ising model in statistical mechanics and statistics, phase transition (see Georgii, 2011), which we explain in Appendix H.1.
@@
The simulation study was run on two servers with Intel processors (Xeon Gold 5220 (2nd gen.) and Xeon Gold 5420+ (4th gen.)). Fitting the model to a single dataset used a single thread. Across all datasets of the study on the 2nd gen. processor, the average elapsed computation time of the aMRF was the fastest, with an average of 6.0 minutes to obtain five thousand draws from the posterior. The two MDGM models had similar average elapsed times, 20.0 minutes for the spanning tree class and 20.8 minutes for the acyclic orientation class. The exact MRF model was nearly eight times slower than the MDGM models, with an average elapsed time of 65.5 minutes. On the 4th gen. processor, the averaged elapsed times were 3.3 minutes for the aMRF, 13.6 for the MDGM-AO and MDGM-ST, and 44.1 minutes for the exact MRF.
@@
The results of the simulation study, the estimates of the expected statistic with a 90% bootstrap confidence interval of the Monte Carlo error, are shown in Figure 1. The bootstrap confidence interval estimates were obtained by taking one thousand samples, of size one hundred and with replacement, of the expected posterior statistics across the datasets, then taking the average of each resample to estimate the sampling distribution. We took the fifth and ninety-fifth quantiles of the estimated sampling distribution to be our lower and upper bounds of the confidence interval.
@@
The three metrics each share different information on model performance. All models perform equally well in terms of accuracy, in both the high ($\lambda=1.39$, 25% missing rate) and low uncertainty ($\lambda=2.3$, 10% missing rate) settings; The expected areal-unit-wise error rate is about the same regardless of prior. The models do not perform equally well in capturing the true spatial dependence of the latent process. In particular, the aMRF model has a higher expected posterior RMSE as the spatial dependence parameter increases. This worse performance for $\beta\geq 0.5$ is more pronounced in the high uncertainty setting. The MDGM-ST consistently performs on par with the MRF in recovering the true spatial dependence. While there is greater uncertainty in the Monte Carlo estimates for the MDGM-AO, this model also performs well and beats the aMRF in the high uncertainty setting when $\beta$ is large. Lastly, the total variation distance gives us a measure of approximation when an MRF is the true data generating distribution. Total variation distance ranges from zero to one. For all three substitute priors, the posterior probabilities over the number of matches with respect to the NUG do differ from that of the posterior for a model with an MRF prior. In both the low and high uncertainty regime, it is clear that the MDGM-ST is a better approximation to the MRF than the MDGM-AO or aMRF priors. While not presented as an approximation, the MDGM-ST turns out to be a better approximation to an MRF than the commonly used aMRF.
Across all metrics, the spanning tree class performs the best. As spanning trees are strongly compatible, conditionally, false dependencies not present in the NUG are not introduced. As a result, spanning trees can come closer to approximating an MRF distribution which perfectly encodes the dependencies of the NUG. In Appendix G.2, we explain that the MDGM cannot represent the exact same set of conditional independence relationships as the NUG due the the implicit indicator variable for the DAG structure; however, conditioning on the DAG, in the hierarchical set up of the model, a collection of strongly compatible DAGs can encapsulate all the conditional independencies represented in the NUG. Notably, the aMRF is a worse representation of the conditional independence relationships of the NUG because it introduces v-structures through its product mixture formulation of the sampling distribution. See Appendix F for more details on how the aMRF is an unnormalized product mixture.